OpenStudy (anonymous):

Solve for n and select the correct answer. 18n + 29 = -9(4 – 2n)

jimthompson5910 (jim_thompson5910):

18n+29=-9(4-2n) 18n+29=-36+18n 18n=-36+18n-29 18n-18n=-36-29 0n=-36-29 0n=-65 0=-65 Since this equation is never true for any value of n, this means that there are no solutions.